3.The ave. weekly income of IT worker is $719. Given the standard deviation is $70. Compute the probability with earnings of 720 dollars/week for a random sample Of 60 IT workers. DO we need to assume the normal distribution? Explain. (Round-up youranswerforz-score to 2 decimal place.)


1
Expert's answer
2021-04-29T17:08:23-0400

"\mu=" $"719"

"\sigma=" $"70"


X=$720,n=60


Yes we need to assume normal distribution since The value of the sample size n is greeter than 30.


using z-score formula for normal sdistribution we have ,

z ="\dfrac{X-\mu}{\sigma}"


z= "\dfrac{(720-719)}{70}=0.01428"


P(z=0.01428)= 0.5057